Problem 41.3 Let X and Y be independent random variables each geometrically distributed with parameter p, i.e. p(1- p otherwise. Find the probability mass function of X +Y
Answer #1

A geometric random variable is the count of Bernouli trial untill a success. we count the probability of obtaining n−1 failures and then 1 success.

PX(n) = p (1−p)n-1   :n=1,2,…

The sum of two such is the count of Bernouli trials untill the 2nd success. the probability of getting 1 success and (n−2) failures, in any arrangement of those (n−1) trials, followed by the second success. is given by:

Px+y(n)=(n−1)(1−p)n−2p2   n=2,3….........

Now

n-1 に!

2-1 〉 , P(X = k)P(Y = n-k) k=1

due to independenc of X and Y

k=1

n-1 n-2 2 n-2 2 に!

which is the required pmf of x+y. where n=2,3,4,........